# Bootstrapped Control Limits for Score-Based Concept Drift Control Charts

### v1

@@ 3 Nested Bootstrap Procedure for Computing the CL
For each $i=1,2,\cdots$, one might consider using the empirical distribution of a naïve version of the bootstrapped $T^{2}$ statistics $\{(\mathbf{z}_{i}^{\,b,j}-\overline{\mathbf{s}}^{\,b})^{\top}(\widehat{\boldsymbol{\Sigma}}^{\,b})^{-1}(\mathbf{z}_{i}^{\,b,j}-\overline{\mathbf{s}}^{\,b}):b=1,2,\cdots,B_{O};j=1,2,\cdots,B_{I}\}$ to approximate the distribution of the $T^{2}$ statistic $T_{n+i}=(\mathbf{z}_{n+i}-\overline{\mathbf{s}})^{\top}\hat{\Sigma}^{-1}(\mathbf{z}_{n+i}-\overline{\mathbf{s}})$ for $\mathcal{S}^{\text{new}}$ and compute its control limit. However, this naïve bootstrap distribution can be severely biased due to the finite sample size of $\mathcal{S}_{OOB}^{b}$, for reasons that become clear later (also see Remark 3.2 below). The following derivations account for this bias and result in control limits that accurately control false-alarm rate.
To derive appropriate control limits, we make the following standard bootstrap assumption.
Assumption 3.1.
(The typical bootstrap assumption) If $\mathcal{D}^{\text{new}}$ is drawn from the same distribution as $\mathcal{D}$, the joint distribution of $(\mathcal{S}_{.368}^{\text{new }},\overline{\mathbf{s}},\widehat{\boldsymbol{\Sigma}},\hat{\boldsymbol{\theta}})$ is the same as the joint distribution of the analogous quantities $\left(\mathcal{S}_{OOB}^{b},\overline{\mathbf{s}}^{b},\widehat{\boldsymbol{\Sigma}}^{b},\hat{\boldsymbol{\theta}}^{b}\right)$ from the bootstrapping procedure, where $\mathcal{S}_{.368}^{\text{new }}$ is any randomly drawn subset of $\mathcal{S}^{\text{new }}$ having the same cardinality $0.368n$ as $\mathcal{S}_{OOB}^{b}$.
The Assumption 3.1, together with (3.1), imply that
$\mathbf{s}_{OOB,i}^{b}|\hat{\boldsymbol{\theta}}^{b}\sim iid\left(\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right),\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right),\quad i=1,2,3,\cdots.$ (3.3)
In the inner loop of the nested bootstrapping procedure of Algorithm 1, $\mathcal{S}_{OOB}^{b}$ serves as the “population” from which the bootstrapped score vectors $\{\mathbf{s}_{i}^{b,j}:i=1,2,3,\cdots\}$ for each inner replicate $j$ are drawn with replacement, having population mean and covariance
$\overline{\mathbf{s}}_{OOB}^{b}=\frac{1}{\lfloor 0.368n\rfloor}\sum_{i=1}^{\lfloor 0.368n\rfloor}\mathbf{s}_{OOB,i}^{b},$ (3.4)
and
$\widehat{\boldsymbol{\Sigma}}_{OOB}^{b}=\frac{1}{\lfloor 0.368n\rfloor}\sum_{i=1}^{\lfloor 0.368n\rfloor}\left(\mathbf{s}_{OOB,i}^{b}-\overline{\mathbf{s}}_{OOB}^{b}\right)\left(\mathbf{s}_{OOB,i}^{b}-\overline{\mathbf{s}}_{OOB}^{b}\right)^{\top},$ (3.5)
where from (3.3),
$\overline{\mathbf{s}}_{OOB}^{b}\lvert\,\hat{\boldsymbol{\theta}}^{b}\sim\left(\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right),\frac{\mathbf{V}(\hat{\boldsymbol{\theta}}^{b})}{\lfloor 0.368n\rfloor}\right),$ (3.6)
and
$\mathbb{E}_{0}\left[\widehat{\boldsymbol{\Sigma}}_{OOB}^{b}\middle|\hat{\boldsymbol{\theta}}^{b}\right]=\frac{\lfloor 0.368n\rfloor-1}{\lfloor 0.368n\rfloor}\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)\cong\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right).$ (3.7)
For each $i=1,2,3,\cdots$, the preceding provides a means to relate the distribution of each $\mathbf{z}_{i}^{b,j}$ computed in the inner bootstrap loop to the distribution of $\mathbf{z}_{n+i}$ for the new data. Since $\{\mathbf{s}_{i}^{b,j}:i=1,2,3,\cdots\}$ is drawn randomly with replacement from $\mathcal{S}_{OOB}^{b}$, it follows from (3.4) - (3.7) that the conditional mean and covariance of $\mathbf{z}_{i}^{b,j}=\lambda\bigl{[}\mathbf{s}_{i}^{\,b,j}+(1-\lambda)\mathbf{s}_{i-1}^{\,b,j}+(1-\lambda)^{2}\mathbf{s}_{i-2}^{\,b,j}+\cdots+(1-\lambda)^{i-1}\mathbf{s}_{1}^{\,b,j}\bigr{]}$ are
$\mathbb{E}_{0}\!\Bigl{[}\mathbf{z}_{i}^{\,b,j}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$ $=\mathbb{E}_{0}\!\Bigl{[}\mathbb{E}_{0}\!\Bigl{\{}\lambda\bigl{[}\mathbf{s}_{i}^{\,b,j}+(1-\lambda)\mathbf{s}_{i-1}^{\,b,j}+(1-\lambda)^{2}\mathbf{s}_{i-2}^{\,b,j}+\cdots+(1-\lambda)^{i-1}\mathbf{s}_{1}^{\,b,j}\bigr{]}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{\}}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$ (3.8)
$=\mathbb{E}_{0}\!\Bigl{[}\lambda\mathbb{E}_{0}\!\Bigl{[}\mathbf{s}_{i}^{\,b,j}\Bigm{|}\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{]}+\lambda(1-\lambda)\mathbb{E}_{0}\!\Bigl{[}\mathbf{s}_{i-1}^{\,b,j}\Bigm{|}\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{]}$
$\qquad\quad+\cdots+\lambda(1-\lambda)^{i-1}\mathbb{E}_{0}\!\Bigl{[}\mathbf{s}_{1}^{\,b,j}\Bigm{|}\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{]}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$
$=\mathbb{E}_{0}\!\Bigl{[}\lambda\sum_{k=0}^{i-1}(1-\lambda)^{k}\,\overline{\mathbf{s}}_{OOB}^{\,b}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}\lambda\sum_{k=0}^{i-1}(1-\lambda)^{k}\,\mathbb{E}_{0}\!\Bigl{[}\overline{\mathbf{s}}_{OOB}^{\,b}\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}=\bigl{[}1-(1-\lambda)^{i}\bigr{]}\,\boldsymbol{\mu}\!\bigl{(}\hat{\boldsymbol{\theta}}^{\,b}\bigr{)}.$
and
$\operatorname{Cov}_{0}\!\Bigl{[}\mathbf{z}_{i}^{\,b,j}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$ $=\mathbb{E}_{0}\!\Bigl{[}\operatorname{Cov}\!\Bigl{(}\lambda\bigl{[}\mathbf{s}_{i}^{\,b,j}+(1-\lambda)\mathbf{s}_{i-1}^{\,b,j}+(1-\lambda)^{2}\mathbf{s}_{i-2}^{\,b,j}+\cdots+(1-\lambda)^{i-1}\mathbf{s}_{1}^{\,b,j}\bigr{]}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{)}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$ (3.9)
$\quad+\operatorname{Cov}_{0}\!\Bigl{[}\mathbb{E}_{0}\!\Bigl{(}\lambda\bigl{[}\mathbf{s}_{i}^{\,b,j}+(1-\lambda)\mathbf{s}_{i-1}^{\,b,j}+(1-\lambda)^{2}\mathbf{s}_{i-2}^{\,b,j}+\cdots+(1-\lambda)^{i-1}\mathbf{s}_{1}^{\,b,j}\bigr{]}\,\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{)}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$
$=\mathbb{E}_{0}\!\Bigl{[}\lambda^{2}\sum_{k=0}^{i-1}(1-\lambda)^{2k}\operatorname{Cov}_{0}\!\Bigl{[}\mathbf{s}_{i-k}^{\,b,j}\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{]}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$
$\qquad\quad+\operatorname{Cov}_{0}\!\Bigl{\{}\lambda\sum_{k=0}^{i-1}(1-\lambda)^{k}\mathbb{E}_{0}\!\Bigl{[}\mathbf{s}_{i-k}^{\,b,j}\Bigm{|}\,\hat{\boldsymbol{\theta}}^{\,b},\mathcal{D}_{\mathrm{OOB}}^{\,b}\Bigr{]}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{\}}$
$=\mathbb{E}_{0}\!\Bigl{[}\lambda^{2}\sum_{k=0}^{i-1}(1-\lambda)^{2k}\widehat{\boldsymbol{\Sigma}}_{OOB}^{b}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}+\operatorname{Cov}_{0}\!\Bigl{[}\lambda\sum_{k=0}^{i-1}(1-\lambda)^{k}\overline{\mathbf{s}}_{OOB}^{b}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}$
$=\mathbb{E}_{0}\!\Bigl{[}\tfrac{\lambda}{2-\lambda}\,\hat{\boldsymbol{\Sigma}}_{\mathrm{OOB}}^{\,b}\bigl{[}1-(1-\lambda)^{2i}\bigr{]}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{]}+\operatorname{Cov}_{0}\!\Bigl{\{}\bigl{[}1-(1-\lambda)^{i}\bigr{]}\,\overline{\mathbf{s}}_{\mathrm{OOB}}^{\,b}\,\Bigr{|}\,\hat{\boldsymbol{\theta}}^{\,b}\Bigr{\}}$
$\cong\tfrac{\lambda}{2-\lambda}\,\bigl{[}1-(1-\lambda)^{2i}\bigr{]}\,\mathbf{V}\!\bigl{(}\hat{\boldsymbol{\theta}}^{\,b}\bigr{)}+\tfrac{1}{0.368\,n}\,\bigl{[}1-(1-\lambda)^{i}\bigr{]}^{2}\,\mathbf{V}\!\bigl{(}\hat{\boldsymbol{\theta}}^{\,b}\bigr{)}$
=$=\Bigl{\{}\tfrac{\lambda}{2-\lambda}\,\bigl{[}1-(1-\lambda)^{2i}\bigr{]}+\tfrac{1}{0.368\,n}\,\bigl{[}1-(1-\lambda)^{i}\bigr{]}^{2}\Bigr{\}}\mathbf{V}\!\bigl{(}\hat{\boldsymbol{\theta}}^{\,b}\bigr{)}.$
The additive term $[1-(1-\lambda)^{i}]^{2}/0.368n$ quantifies the extra variability introduced by estimating the mean score from an out-of-bag sample that contains on average $(1-0.632)n=0.368n$ points rather than the full training size. Including this 632-style correction inflates the conditional variance to the appropriate level and is essential for the resulting CL to achieve the target false-alarm rate (also see Remark 3.2 below).
From (3.8) and (3.9), the unconditional mean and covariance of $\mathbf{z}_{i}^{b,j}$ are
$\mathbb{E}_{0}\left[\mathbf{z}_{i}^{b,j}\right]=\mathbb{E}_{0}\left[\mathbb{E}_{0}\left[\mathbf{z}_{i}^{b,j}\middle|\hat{\boldsymbol{\theta}}^{b}\right]\right]=\left[1-(1-\lambda)^{i}\right]\mathbb{E}_{0}\left[\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right],$ (3.10)
and
$\operatorname{Cov}_{0}\left[\mathbf{z}_{i}^{b,j}\right]$ $=\mathbb{E}_{0}\left[\operatorname{Cov}_{0}\left[\mathbf{z}_{i}^{b,j}\middle|\hat{\boldsymbol{\theta}}^{b}\right]\right]+\operatorname{Cov}_{0}\left[\mathbb{E}_{0}\left[\mathbf{z}_{i}^{b,j}\middle|\hat{\boldsymbol{\theta}}^{b}\right]\right]$ (3.11)
=$=\left\{\frac{\lambda}{2-\lambda}[1-(1-\lambda)^{2i}]+\frac{1}{0.368n}[1-(1-\lambda)^{i}]^{2}\right\}\mathbb{E}_{0}\left[\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right]+\bigl{[}1-(1-\lambda)^{i}\bigr{]}^{2}\,\operatorname{Cov}_{0}\!\bigl{[}\boldsymbol{\mu}(\hat{\boldsymbol{\theta}}^{\,b})\bigr{]}.$
To approximate the terms on the right hand sides of (3.10) and (3.11) and relate them to (3.2), we denote by $(\mathbf{X},Y)$ a predictor-response pair randomly drawn from the population $\mathbb{P}_{0}(Y,\mathbf{X})$, or equivalently from $\mathcal{D}_{OOB}^{b}$, and we consider the first-order Taylor approximations of $\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right)$ and $\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)$ about $\hat{\boldsymbol{\theta}}^{b}\cong\boldsymbol{\theta}^{0}$, where $\boldsymbol{\theta}^{0}$ denotes the true values of the parameters. Assume that the functions $\mathbf{\mu}(\boldsymbol{\theta})$ and $\mathbf{V}(\boldsymbol{\theta})$ are continuously differentiable with respect to $\boldsymbol{\theta}^{0}$ in a neighborhood of $\boldsymbol{\theta}^{0}$. Then, it holds that
$\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right)$ $\cong\mathbf{\mu}\left(\boldsymbol{\theta}^{0}\right)+\left.\nabla_{\boldsymbol{\theta}^{\top}}\mathbf{\mu}(\boldsymbol{\theta})\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)$ (3.12)
$=\mathbb{E}_{0}\left[\left.\nabla_{\boldsymbol{\theta}}\log\mathbb{P}\left(Y\middle|\mathbf{X};\boldsymbol{\theta}\right)\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\right]+\left.\nabla_{\boldsymbol{\theta}^{\top}}\mathbb{E}_{0}\left[\nabla_{\boldsymbol{\theta}}\log\mathbb{P}\left(Y\middle|\mathbf{X};\boldsymbol{\theta}\right)\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)$
$=\mathbf{0}+\left.\mathbb{E}_{0}\left[\nabla^{2}_{\boldsymbol{\theta}}\log\mathbb{P}\left(Y\middle|\mathbf{X};\boldsymbol{\theta}\right)\right]\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)=-\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)$ , UNKNOWN
where we assume sufficient regularity conditions on the log-likelihood to allow the expectation and gradient operators to be exchanged, $\mathbf{I}(\boldsymbol{\theta}^{0})$ denotes the expected Fisher information matrix, and we have used the standard result that, at the true parameters $\boldsymbol{\theta}^{0}$, the score function is zero-mean. Similarly,
$\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)\cong\mathbf{V}\left(\boldsymbol{\theta}^{0}\right)+\left.\nabla_{\boldsymbol{\theta}^{\top}}\mathbf{V}(\boldsymbol{\theta})\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)=\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)+\nabla_{\boldsymbol{\theta}^{\top}}\mathbf{V}(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right),$ (3.13)
where we have used the standard result that the covariance matrix of the score vector at $\boldsymbol{\theta}^{0}$ is $\mathbf{I}(\boldsymbol{\theta}^{0})$. Using the standard asymptotic result $\hat{\boldsymbol{\theta}}{\sim}\mathcal{N}(\boldsymbol{\theta}^{0},\mathbf{I}^{-1}(\boldsymbol{\theta}^{0})/n)$ (which by Assumption 3.1 is also the asymptotic distribution of $\hat{\boldsymbol{\theta}}^{\,b}$) with (3.12) and (3.13) gives
$\mathbb{E}_{0}\left[\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right]$ $\cong\mathbb{E}_{0}\left[-\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)\right]\cong\mathbf{0},$ (3.14)
$\operatorname{Cov}_{0}\left[\mathbf{\mu}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right]\cong\operatorname{Cov}_{0}\left[-\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)\right]\cong\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)\frac{\mathbf{I}^{-1}(\boldsymbol{\theta}^{0})}{n}\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)=\frac{\mathbf{I}(\boldsymbol{\theta}^{0})}{n},$ (3.15)
and
$\mathbb{E}_{0}\left[\mathbf{V}\left(\hat{\boldsymbol{\theta}}^{b}\right)\right]\cong\mathbb{E}_{0}\left[\mathbf{I}\left(\boldsymbol{\theta}^{0}\right)+\left.\frac{\partial\mathbf{V}(\boldsymbol{\theta}^{0})}{\partial\boldsymbol{\theta}^{\top}}\right|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{0}}\left(\hat{\boldsymbol{\theta}}^{b}-\boldsymbol{\theta}^{0}\right)\right]=\mathbf{I}\left(\boldsymbol{\theta}^{0}\right).$ (3.16)
Substituting (3.14)-(3.16) into (3.10) and (3.11) gives
$\mathbb{E}_{0}\left[\mathbf{z}_{i}^{b,j}\right]\cong\mathbf{0},\quad\operatorname{Cov}_{0}\left[\mathbf{z}_{i}^{b,j}\right]=\left\{\frac{\lambda}{2-\lambda}[1-(1-\lambda)^{2i}]+\frac{3.72}{n}[1-(1-\lambda)^{i}]^{2}\right\}\mathbf{I}\left(\boldsymbol{\theta}^{0}\right).$ (3.17)
Using equations (3.2) and repeating the derivations of (3.12)-(3.16) for $\hat{\boldsymbol{\theta}}$ instead of for $\hat{\boldsymbol{\theta}}^{b}$ gives the analogous results
$\mathbb{E}_{0}\left[\mathbf{z}_{i}\right]\cong\mathbf{0},\quad\operatorname{Cov}_{0}\left[\mathbf{z}_{i}\right]\cong\left\{\frac{\lambda}{2-\lambda}[1-(1-\lambda)^{2i}]+\frac{1}{n}[1-(1-\lambda)^{i}]^{2}\right\}\mathbf{I}\left(\boldsymbol{\theta}^{0}\right).$ (3.18)
Comparing (3.17) with (3.18), we see that $\mathbf{z}_{i}^{b,j}$ and $\mathbf{z}_{i}$ have the same mean $\mathbf{0}$ and covariance matrices that are scalar multiples of $\mathbf{I}(\boldsymbol{\theta}^{0})$. Relative to $\mathbf{z}_{i}$, the covariance matrix of $\mathbf{z}_{i}^{b,j}$ is inflated by the factor
$k\left(\lambda,i,n\right):=\frac{\frac{\lambda}{2-\lambda}[1-(1-\lambda)^{2i}]+\frac{3.72}{n}[1-(1-\lambda)^{i}]^{2}}{\frac{\lambda}{2-\lambda}[1-(1-\lambda)^{2i}]+\frac{1}{n}[1-(1-\lambda)^{i}]^{2}}.$ (3.19)
This provides the basis for taking the CL function $CL_{i}$ in Algorithm 1 to be the upper $\alpha$ quantile of the quantities $\{T_{i}^{b,j}:b=1,2,\cdots,B_{0};j=1,2,\cdots,B_{I}\}$, which serves as an approximation to the upper $\alpha$ quantile of the distribution of $T_{n+i}$, where
$T_{i}^{b,j}\equiv\left(\frac{\mathbf{z}_{i}^{b,j}}{\sqrt{k(\lambda,i,n)}}-\overline{\mathbf{s}}^{b}\right)^{\top}\left(\widehat{\boldsymbol{\Sigma}}^{b}\right)^{-1}\left(\frac{\mathbf{z}_{i}^{b,j}}{\sqrt{k(\lambda,i,n)}}-\overline{\mathbf{s}}^{b}\right),$ (3.20)
and
$T_{n+i}\equiv\left(\mathbf{z}_{n+i}-\overline{\mathbf{s}}\right)^{\top}\widehat{\boldsymbol{\Sigma}}^{-1}\left(\mathbf{z}_{n+i}-\overline{\mathbf{s}}\right).$ (3.21)

### v2

@@ 3 Nested Bootstrap Procedure for Computing the CL
For each $i=1,2,\cdots$, one might consider using the empirical distribution of a naïve version of the bootstrapped $T^{2}$ statistics $\{(\mathbf{z}_{i}^{\,b,j}-\overline{\mathbf{s}}^{\,b})^{\top}(\widehat{\boldsymbol{\Sigma}}^{\,b})^{-1}(\mathbf{z}_{i}^{\,b,j}-\overline{\mathbf{s}}^{\,b}):b=1,2,\cdots,B_{O};j=1,2,\cdots,B_{I}\}$ to approximate the distribution of the $T^{2}$ statistic $T_{n+i}=(\mathbf{z}_{n+i}-\overline{\mathbf{s}})^{\top}\hat{\Sigma}^{-1}(\mathbf{z}_{n+i}-\overline{\mathbf{s}})$ for $\mathcal{S}^{\text{new}}$ and compute its CL. However, this naïve bootstrap distribution can be severely biased due to the finite sample size of $\mathcal{S}_{OOB}^{b}$, for reasons that become clear later (also see Remark 3.3 below). The following derivations account for this bias and result in CLs that accurately control false-alarm rate.
3.1 Covariance Inflation and Control Limit Construction
[…]
The main result of these derivations is that the bootstrap MEWMA statistic is a stochastically inflated version of the true MEWMA statistic. This motivates the scaling factor $k(\lambda,i,n)$ presented below.
Theorem 3.1 (Covariance inflation factor).
[…]
$k(\lambda,i,n)=\frac{\displaystyle\frac{\lambda}{2-\lambda}\bigl[1-(1-\lambda)^{2i}\bigr]+\frac{3.72}{n}\bigl[1-(1-\lambda)^{i}\bigr]^{2}}{\displaystyle\frac{\lambda}{2-\lambda}\bigl[1-(1-\lambda)^{2i}\bigr]+\frac{1}{n}\bigl[1-(1-\lambda)^{i}\bigr]^{2}}$}, (3.3)
in the sense that
$\operatorname{Cov}_{0}\!\,\bigl[\mathbf{z}^{b,j}_{i}\bigr]\cong k(\lambda,i,n)\,\operatorname{Cov}_{0}\!\,\bigl[\mathbf{z}_{i}\bigr].$} (3.4)
Theorem 3.1 implies that dividing the bootstrap MEWMA statistic by $\sqrt{k(\lambda,i,n)}$ produces a variance-corrected quantity whose covariance matches that of the true in-control MEWMA statistic. This leads to the following procedure for determining the Control Limits.
Corollary 3.2 (Variance-corrected bootstrap $T^{2}$ and control limits).
Let $k(\lambda,i,n)$ be as in (3.3) and define the variance-corrected bootstrap MEWMA statistic $\tilde{\mathbf{z}}^{\,b,j}_{i}=\mathbf{z}^{b,j}_{i}/\sqrt{k(\lambda,i,n)}.$ For each monitoring time $i\geq 1$, define the bootstrap Hotelling $T^{2}$ statistics
$T^{b,j}_{i}=\bigl(\tilde{\mathbf{z}}^{\,b,j}_{i}-\bar{\mathbf{s}}^{\,b}\bigr)^{\!\top}\bigl(\hat{\boldsymbol{\Sigma}}^{\,b}\bigr)^{-1}\bigl(\tilde{\mathbf{z}}^{\,b,j}_{i}-\bar{\mathbf{s}}^{\,b}\bigr),$} (3.5)
and take the CL $\mathrm{CL}_{i}$ to be the empirical upper $(1-\alpha)$ quantile of $\{T^{b,j}_{i}:b=1,\ldots,B_{O},\;j=1,\ldots,B_{I}\}$. Then the resulting MEWMA chart applied to
$T_{n+i}=\bigl(\mathbf{z}_{n+i}-\bar{\mathbf{s}}\bigr)^{\!\top}\hat{\boldsymbol{\Sigma}}^{-1}\bigl(\mathbf{z}_{n+i}-\bar{\mathbf{s}}\bigr)$} (3.6)
has point-wise false-alarm probability approximately $\alpha$ at each time $i$ under $\mathbb{P}_{0}$.
After the CLs are obtained, monitoring proceeds in the usual way. For each new incoming observation $(\mathbf{X}_{n+i},Y_{n+i})$, compute its score vector, update the MEWMA statistic $\mathbf{z}_{n+i}$, and compare the corresponding $T_{n+i}$ value to $\mathrm{CL}_{i}$. A signal is issued at the first time $i$ for which $T_{n+i}>\mathrm{CL}_{i}$.
3.2 Stabilization and Detectability of the Control Limit
Throughout the remainder of this section, recall that $\mathcal{D}$ denotes the fixed Phase I training sample. Let $T_{i}^{*}$ denote a generic bootstrap replicate of the monitoring statistic at time $i$, constructed according to the resampling scheme in Algorithm 1 and Corollary 3.2. Conditional on $\mathcal{D}$, let $F_{i}(\cdot\mid\mathcal{D})$ denote the distribution function of $T_{i}^{*}$, which accounts for both Phase I estimation uncertainty and Phase II variability, and let $q_{i,\alpha}(\mathcal{D})$ be its $(1-\alpha)$ quantile. To describe the steady-state behavior, define $T_{\infty}^{*}$ to be the stationary analogue of $T_{i}^{*}$ obtained by applying the same recursion to a stationary (two-sided) bootstrap score sequence. Let $F_{\infty}(\cdot\mid\mathcal{D})$ denote the conditional distribution of $T_{\infty}^{*}$ given $\mathcal{D}$, and let $q_{\infty,\alpha}(\mathcal{D})$ denote its $(1-\alpha)$ quantile.
Because the MEWMA statistic is initialized at zero, the bootstrap control limits $\mathrm{CL}_{i}$ defined in Corollary 3.2 are time-varying at early monitoring times. In long monitoring runs, however, recomputing the control limit indefinitely is computationally inefficient. The following result shows that the bootstrap control limit stabilizes geometrically fast as the monitoring time increases.
Theorem 3.3 (Stabilization of the bootstrap control limit).
Assume Assumptions A.1A.4 hold and that there is no concept drift. In addition, assume that the limiting bootstrap distribution $F_{\infty}(\cdot\mid\mathcal{D})$ of $T_{\infty}^{*}$ admits a density $f_{\infty}(\cdot\mid\mathcal{D})$ such that: (i) $f_{\infty}(q_{\infty,\alpha}(\mathcal{D})\mid\mathcal{D})\geq c_{0}(\mathcal{D})>0$, for some constant $c_{0}(\mathcal{D})$, and (ii) $f_{\infty}(\cdot\mid\mathcal{D})$ is bounded in a neighborhood of $q_{\infty,\alpha}(\mathcal{D})$. Then $q_{\infty,\alpha}(\mathcal{D})$ is finite and
$\bigl|q_{i,\alpha}(\mathcal{D})-q_{\infty,\alpha}(\mathcal{D})\bigr|\;\leq\;C(\mathcal{D})\,(1-\lambda)^{i}\quad\text{for all }i\geq 1,$
where $C(\mathcal{D})<\infty$ depends only on the Phase I training sample $\mathcal{D}$.